# Optimizing Attention with Mirror Descent: Generalized Max-Margin Token Selection

### v3

@@ G.2 $\ell_{p}$-AttGD Experiment
The dataset $(X_{i},Y_{i},z_{i})_{i=1}^{n}$ is generated randomly: $X_{i}$ and $z_{i}$ are sampled from the unit sphere, and $Y_{i}$ is uniformly sampled from {±$\{\pm 1\}$. Additionally, $v$ is randomly selected from the unit sphere. We use $n=6$ samples, $T=8$ tokens per sample, and $d=10$ dimensions per token, fulfilling the overparameterized condition for the $\ell_{p}$-AttSVM problem to be almost always feasible.
@@
The model is trained with parameters initialized near the origin, using Algorithms $\ell_{p}$-AttGD with $p=1.75,2$, and $3$, and a learning rate of $0.1$. Training lasted for $1,500$ epochs for $p=1.75$, $2,000$ epochs for $p=2$, and $20,000$ epochs for $p=3$. Gradients are normalized to accelerate convergence without altering results significantly. We refer to the parameter histories as the $\ell_{1.75},\ell_{2},$ and $\ell_{3}$ optimization paths. We compute the chosen tokens $(\alpha_{i})_{i=1}^{n}$ for the ($\ell_{p}$-AttSVM) problem by selecting the token with the highest softmax probability for each sample. This process is repeated for $p=1.75$, $2$, and $3$.
@@ G.4 Architecture Details for Stanford Large Movie Review Classification
The embedding layer has trainable token embeddings $E$ and position encodings $P$. The model’s vocabulary size is $30{,}522$, with a maximum token length of $512$ and an embedding dimension of $384$. Hence, $E\in\mathbb{R}^{30522\times 384}$ and $P\in\mathbb{R}^{512\times 384}$. If a token $t$ is in position $i$, its embedding is $X_{i}=E_{t}+P_{i}$, where $E_{t}$ and $P_{i}$ denote the $t^{\text{th}}$ and $i^{\text{th}}$ rows of $E$ and $P$, respectively.
@@ G.5 Additional Experiment with Adam
The ViT architecture that we trained to compare $\ell_{1.1}$ and Adam used a patch size of $4$ on the input image, $512$ dimensional token feature, $6$ layers of attention blocks with $8$ attention heads per attention layer, and a one-layer linear classification layer for the final prediction head on the final [CLS] patch token feature.
The embedding layer of the architecture follows the work of Dosovitskiy et al. (2020), where the embedding layer learns the [CLS] token embedding, a linear map for embedding each image patch, and a positional embedding for each possible position on the image. The details of how the attention layers are implemented are similar to that of the architecture used for the Stanford Large Movie dataset, just that the MLP sublayer is now a two-layer GeLU network, and that we apply the layer normalization before the multihead attention and the MLP instead of after the residual connection. Furthermore, we apply a dropout of $0.1$ when training the network.
@@ G.6 Addendum to the Attention Map Results
Figure 13: The full attention map table that shows that $\ell_{1.1}$-MD provides strictly more attention to the pivotal token compared to $\ell_{2}$-MD, or equivalently GD, for $22$ of the sample sentences. Out of the other $18$ sentences, for $16$ of which, GD strictly outperforms $\ell_{1.1}$-MD, while for the other $2$, the two algorithms are equally as good.

### v5

@@ G.2 $\ell_{p}$-AttGD Experiment
The dataset $(X_{i},Y_{i},z_{i})_{i=1}^{n}$ is generated randomly: $X_{i}$ and $z_{i}$ are sampled from the unit sphere, and $Y_{i}$ is uniformly sampled from $\{\pm 1\}$. Additionally, $v$ is randomly selected from the unit sphere. We use $n=6$ samples, $T=8$ tokens per sample, and $d=10$ dimensions per token, fulfilling the overparameterized condition for the $\ell_{p}$-AttSVM problem to be almost always feasible.
@@
The model is trained with parameters initialized near the origin, using $\ell_{p}$-AttGD with $p=1.75,2$, and $3$, and a learning rate of $0.1$. Training lasted for $1,500$ epochs for $p=1.75$, $2,000$ epochs for $p=2$, and $20,000$ epochs for $p=3$. Gradients are normalized to accelerate convergence without altering results significantly. We refer to the parameter histories as the $\ell_{1.75},\ell_{2},$ and $\ell_{3}$ optimization paths. We compute the chosen tokens $(\alpha_{i})_{i=1}^{n}$ for the ($\ell_{p}$-AttSVM) problem by selecting the token with the highest softmax probability for each sample. This process is repeated for $p=1.75$, 22, and $3$.
@@ G.4 Architecture Details for Stanford Large Movie Review Classification
The embedding layer has trainable token embeddings $E$ and position encodings $P$. The model’s vocabulary size is 30,52230{,}522, with a maximum token length of $512$ and an embedding dimension of $384$. Hence, $E\in\mathbb{R}^{30522\times 384}$ and $P\in\mathbb{R}^{512\times 384}$. If a token $t$ is in position $i$, its embedding is $X_{i}=E_{t}+P_{i}$, where $E_{t}$ and $P_{i}$ denote the $t^{\text{th}}$ and $i^{\text{th}}$ rows of $E$ and $P$, respectively.
@@ G.5 Additional Experiment with Adam
The ViT architecture that we trained to compare $\ell_{1.1}$MD and Adam used a patch size of $4$ on the input image, $512$ dimensional token feature, $6$ layers of attention blocks with $8$ attention heads per attention layer, and a one-layer linear classification layer for the final prediction head on the final [CLS] patch token feature.
The embedding layer of the architecture follows the work of Dosovitskiy et al. (2021a), where the embedding layer learns the [CLS] token embedding, a linear map for embedding each image patch, and a positional embedding for each possible position on the image. The details of how the attention layers are implemented are similar to that of the architecture used for the Stanford Large Movie dataset, just that the MLP sublayer is now a two-layer GeLU network, and that we apply the layer normalization before the multihead attention and the MLP instead of after the residual connection. Furthermore, we apply a dropout of $0.1$ when training the network.
@@ G.6 Addendum to the Attention Map Results
Figure 14: The full attention map table that shows that $\ell_{1.1}$-MD provides strictly more attention to the pivotal token compared to $\ell_{2}$-MD, or equivalently GD, for $22$ of the sample sentences. Out of the other $18$ sentences, for $16$ of which, GD strictly outperforms $\ell_{1.1}$-MD, while for the other 22, the two algorithms are equally as good.
G.7 Additional Results on Attention Weight Sparsity and Runtime Performance
Figure 15: Histogram of the absolute values of the entries in the $W_{K},W_{Q}$, and $W_{V}$ attention weight matrices of the resulting models trained by $\ell_{1.75},\ell_{2}$, and $\ell_{3}$MD on CIFAR-10.
Figure 16: Histogram of the absolute values of the entries in the $W_{K},W_{Q}$, and $W_{V}$ attention weight matrices of the resulting models trained by $\ell_{1.75},\ell_{2}$, and $\ell_{3}$MD on CIFAR-100.
Similar to the histogram plots for $\ell_{1.1}$MD and Adam in Figures 11 and 12 of Section 4.2.2, we plot the histograms for $\ell_{1.75},\ell_{2},$ and $\ell_{3}$MD below. As we can see in Figure 15 and 16, the $\ell_{1.75}$MD histogram shows a sparser weight distribution than that of Adam and $\ell_{2}$MD, while the $\ell_{3}$MD histogram shows a denser distribution.
(a) CIFAR-10 Training Time in Hour
(b) CIFAR-100 Training Time in Hour
Figure 17: The cumulative time taken to train ViT on CIFAR-10 (left) and CIFAR-100 (right) using $\ell_{p}$-MD for $p=1.1,1.75,2,$ and 33 and Adam algorithms across the 2000 epochs. On the left subgraph, it took approximately 29 hours for each of the five optimizers to fully train the model on the CIFAR-10 dataset, and on the right subgraph, it took approximately 25 hours for each of the five optimizers to fully train the model on the CIFAR-100 dataset.
Next, we also take the cumulative time taken to perform each training steps for the experiments in Section 4.2.2. Figure 17 shows that for training CIFAR-10 and CIFAR-100, the differences in the training time between the five different training algorithms are very marginal. Therefore, our algorithms do not have any significant slowdown in the training speed.